# Look before Transcription: End-to-End SlideASR with Visually-Anchored Policy Optimization

### v1

@@ Abstract
Automatic speech recognition (ASR) systems often struggle with domain-specific terminology, especially in specialized settings such as academic lectures. To address this, we define the SlideASR task, which leverages the rich visual information from presentation slides to improve transcription accuracy. Existing pipeline methods for this task tend to be complex and underperform. Although omni-modal large language models (OLLMs) provide a promising end-to-end framework, they frequently fail in practice by degenerating into simple optical character recognition (OCR) systems. To overcome this, we propose Visually-Anchored Policy Optimization (VAPO), a novel post-training method designed to control the model’s reasoning process. Drawing on the Chain-of-Thought reasoning paradigm, VAPO enforces a structured "Look before Transcription" procedure using a <think><answer> format. Specifically, the model first performs OCR on the slide content within the think step, then generates the transcription by referencing this recognized visual information in the answer step. This reasoning process is optimized via reinforcement learning with four distinct rewards targeting format compliance, OCR accuracy, ASR quality, and visual anchoring consistency. To support further research, we construct SlideASR-Bench, a new entity-rich benchmark consisting of a synthetic dataset for training and testing, and a challenging real-world set for evaluation. Extensive experiments demonstrate that VAPO significantly improves recognition of domain-specific terms, establishing an effective end-to-end paradigm for SlideASR.
Look before Transcription: End-to-End SlideASR with
Visually-Anchored Policy Optimization
[…]
@@ 1 Introduction
Current end-to-end automatic speech recognition (ASR) models, such as Whisper Radford et al. (2023), have demonstrated impressive performance in transcribing common words. However, recognition performance often deteriorates significantly in specialized domains, such as academic lectures, technical presentations, or medical seminars. Previous works have improved ASR accuracy by incorporating lip movement information from the speaker Afouras et al. (2022); Ma et al. (2021, 2023); Shi et al. (2022). However, in addition to facial information, there is multimodal textual information on the slide that is closely related to the current speech of the speaker Wang et al. (2024b); Zhao et al. (2025), which lipreading-based approaches cannot utilize. For clarity, we refer to the task of improving ASR accuracy by incorporating visual cues from presentation slides as SlideASR.
Figure 1: We compare the outputs of Qwen2.5-Omni-7B, Qwen3-Omni-30B-A3B (with OCR text as context), and our VAPO-7B on a real Chinese medical report sample. Red text indicates incorrectly transcribed named entities.
Currently, the dominant strategy for SlideASR task is the pipeline paradigm. Specifically, these methods can be divided into two categories. The first is post-processing correction Trinh et al. (2025), which utilizes large language models (LLMs) for text-level refinement after independent ASR and optical character recognition (OCR). The second is contextual enhancement Wang et al. (2024b, a); Yu et al. (2024); Yang et al. (2024), where OCR-extracted text is injected as context into the ASR model. The post-processing correction methods involve multiple modules, resulting in a complex workflow and relatively high latency. In contrast, for contextual enhancement methods, we observed that when OCR text is provided to large audio language models (LALMs) Chu et al. (2024); Dinkel et al. (2025), the models often tend to repeat the context rather than effectively assisting the recognition process.
Beyond the pipeline paradigm, we argue that the recently emerging omni-modal large language models (OLLMs) Xu et al. (2025a, b); Yao et al. (2024); Li et al. (2025a) are inherently well-suited for the SlideASR task. OLLMs are capable of simultaneously processing textual, visual, and auditory modalities, and performing tasks based on human instructions. Therefore, they are theoretically able to accomplish the SlideASR task in an end-to-end manner. However, we identify two key issues with existing OLLMs when applied to the SlideASR task. First, their behavior is often uncontrollable. When given an audio clip and the corresponding slide image, the models do not consistently perform accurate transcription. In some cases, they simply copy text from the slide, effectively acting as OCR systems instead of transcribing speech. Second, even when ASR is successful, the model often fails to leverage entity information from the slide. Entities that are clearly shown in the image are sometimes transcribed incorrectly, suggesting weak integration of visual context. These issues highlight that the implicit and opaque reasoning processes of OLLMs are unreliable for complex multimodal tasks such as SlideASR. Inspired by recent advances in LLMs, where explicit Chain-of-Thought (CoT) reasoning Guo et al. (2025); Jaech et al. (2024); Shao et al. (2024b) has been shown to improve performance on complex problems, we argue that incorporating structured reasoning into the SlideASR task can help constrain model behavior and address the challenges of multimodal fusion.
To this end, we propose Visually-Anchored Policy Optimization (VAPO), a post-training method tailored to enhance the performance of OLLMs on the SlideASR task. The core idea of VAPO is to replace the model’s implicit and uncontrolled reasoning with an explicit, structured process. Specifically, we enforce the model to generate outputs within a structured think-answer format, compelling it to follow a "Look before Transcription" procedure. Within the <think> block, the model is required to first perform OCR to identify textual information from the slides. Subsequently, in the <answer> block, it must generate the final transcription by referring to the content in <think>, particularly using the entities identified from slides as anchors. Our training is guided by four distinct reward functions: a) Format Reward to enforce compliance with the structured output format; b) OCR Reward to promote precise extraction of textual information from slides; c) ASR Reward aimed at enhancing the overall ASR performance; and d) Visual Anchoring Reward that encourages the model to utilize the entities identified within the <think> block as references during the generation of the final transcription. Fig 1 shows an example illustrating the performance of our method. In a challenging real-world example, both a naive OLLM and a pipeline-based method fail to correctly transcribe all domain-specific entities. In contrast, our VAPO-7B model first identifies all entities in the <think> block and subsequently generates a accurate transcription, showcasing the practical advantage of our approach.
Existing SlideASR datasets, such as SlideSpeech Wang et al. (2024b) and ChineseLips Zhao et al. (2025) lack a sufficient number of specialized named entities. To address this, we constructed SlideASR-Bench, a benchmark for the entity-rich SlideASR task, which comprises two subsets: SlideASR-S and SlideASR-R. SlideASR-S is built by synthesizing slides based on entity information from the ContextASR-Bench Wang et al. (2025) dataset. To evaluate model performance in real-world scenarios, we additionally curated SlideASR-R by manually collecting 60 challenging samples from authentic professional reports across four domains: Chemistry, Medicine, Biology, and Artificial Intelligence. See Sec 5 for details.
@@
We conduct experiments on SlideSpeech Wang et al. (2024b), ChineseLips Zhao et al. (2025) and SlideASR-Bench. Results show that our proposed approach outperforms the state-of-the-art models, e.g., Qwen3-Omni-30B-A3B Xu et al. (2025b), particularly on entity-related metrics, while maintaining accuracy on non-entity text. Ablation studies show that each reward function in VAPO is essential for achieving optimal performanc. Attention visualization shows that the VAPO model follows the "Look before Transcription" procedure. The contributions can be summarized as follows:
  • To the best of our knowledge, this work is the first to identify and analyze the limitations of OLLMs when applied to the SlideASR task.
  • We introduce VAPO, a novel post-training method designed to improve the performance of OLLMs in the SlideASR task by enforcing a structured "Look before Transcription" reasoning process.
  • We construct SlideASR-Bench, which consists of two dedicated datasets, SlideASR-S and SlideASR-R, to address the scarcity of domain-specific entities and provide robust benchmarks for the SlideASR task. The data and code will be released.
2 Related Works
Contextual ASR. The objective of contextual ASR is to incorporate contextual information, including domain labels, entity lists, and conversational history, into the speech recognition system in order to improve the recognition accuracy of named entities, and domain-specific terminology Bai et al. (2024); Xiao et al. (2025); Zhou and Li (2025). Besides textual information, researchers have focused on leveraging visual information to enhance the performance of ASR models. For example, integrating lip movement information during the recognition process Ma et al. (2023); Rouditchenko et al. (2024); Shi et al. (2022). This study focuses on the SlideASR task Zhao et al. (2025); Wang et al. (2024b, a), which involves utilizing slide content as contextual information to support the model, given that slides in presentation scenarios generally contain information closely related to the spoken content. Most existing methods for SlideASR are based on the pipeline paradigm Wang et al. (2024a); Zhao et al. (2025); Wang et al. (2024b); Yu et al. (2024); Yang et al. (2024), which results in relatively complex systems. The objective of this study is to accomplish the task using an end-to-end approach.
Omni-modal Large Language Models. Recently, OLLMs Fu et al. (2025); Yao et al. (2024); Xu et al. (2025a); Li et al. (2025a); Hu et al. (2025); Li et al. (2025b) have emerged, integrating vision, audio, and text by aligning their encoders during training for end-to-end processing. Models such as MiniCPM-o Yao et al. (2024) and Qwen2.5-Omni Xu et al. (2025a) have demonstrated strong multimodal performance. Benefiting from the unified modeling capability across visual and audio modalities, they are expected to solve the SlideASR task in an end-to-end manner. However, in practical scenarios, the models exhibit unstable behavior, for instance, they sometimes reproduce the textual content from the slides instead of generating the expected speech transcription.
Chain-of-Thought Reasoning. CoT reasoning is a breakthrough approach that enhances the complex reasoning capabilities of LLMs. Recent works Jaech et al. (2024); Guo et al. (2025) have shown that by using reinforcement learning algorithms Shao et al. (2024b); Rafailov et al. (2023); Schulman et al. (2017) to encourage models to generate intermediate reasoning steps before producing the final answer, performance on tasks involving arithmetic, commonsense, and symbolic reasoning can be significantly enhanced. This paradigm has also been extended to the multimodal domain Shao et al. (2024a); Xu et al. (2024); Ma et al. (2025); Diao et al. (2025), demonstrating the general effectiveness of making reasoning processes explicit.
3 The Failure of OLLMs in SlideASR
3.1 Problem Formulation
Figure 2: Comparison of OLLM outputs with and without slide context.
We formally define the SlideASR task as follows. Given a speech signal $A$ and a slide image $I$ synchronized with the speech, our goal is to train a model $f_{\theta}$ that takes $A$ and $I$ as joint inputs and generates the most likely corresponding text transcription $Y=\{y_{1},y_{2},…,y_{n}\}$, where $y_{n}$ is the $n$-th token in the text sequence. The objective of this task can be expressed as maximizing the conditional probability $P(Y|A,I;\theta)$, where $\theta$ represents the learnable parameters of the model.
$\theta^{*}=\arg\max_{\theta}P(Y|A,I;\theta)$ (1)
In our work, $f_{\theta}$ is an OLLM is that takes both image and audio as input and generates the final text $Y$ in an auto-regressive manner.
3.2 Observations of OLLMs on SlideASR
Although OLLMs can process both auditory and visual modalities simultaneously, we observe that their behavior on the SlideASR task is unstable. Fig 2 shows an example. We selected a sample from SlideSpeech Wang et al. (2024b) and had Qwen2.5-Omni-7B transcribe the speech under two conditions: audio-only and audio with slide image. When the model receives only audio, it produces correct ASR results. However, when both audio and the corresponding slide image are provided, the model ignores the audio signal and instead outputs the text content from the slide image, behaving like a OCR system.
To quantitatively demonstrate the instability of OLLMs’ behavior, we calculated the proportion of OCR-like behavior exhibited by four models Qwen2.5-Omni-7B Xu et al. (2025a), Qwen2.5-Omni-3B Xu et al. (2025a), MiniCPM-o-2.6 Yao et al. (2024), and Megrez-Omni Li et al. (2025a) on the SlideSpeech dataset. The judgment process involves three steps. 1) Identify Common Vocabulary: We first identify the words common to both the ground truth speech transcription and the slide text, denoted as $V_{\text{common}}$. 2) Isolate Slide-Only Vocabulary: We then create a set of slide-only words, $V_{\text{slide\_only}}$, by removing the common words from the slide’s full vocabulary. These words are only accessible to the model via OCR. 3) Detect OCR Behavior: Finally, we check for any intersection between the model’s output and $V_{\text{slide\_only}}$. If such an intersection exists, it indicates that the model has exhibited OCR behavior.
Table 1: Percentage of samples with OCR behavior on SlideSpeech dataset.
Dev set Test set
@@ 3.2 Observations of OLLMs on SlideASR
The results are shown in Table 1. All models exhibited a significant proportion of OCR behavior on the dataset, generating words present on the slide but absent in the audio. This fundamental and widespread failure demonstrates that simply prompting OLLMs is insufficient and that a new, more structured approach is required to control their reasoning process.
4 Visually-Anchored Policy Optimization
To address the issue of uncontrollable behavior in OLLMs for the SlideASR task, we propose Visually-Anchored Policy Optimization (VAPO). The core idea of VAPO is to transform the model from an unreliable "black-box" into a controllable one that follows a structured and effective reasoning path. This is achieved through explicit structured reasoning and multi-objective reward-based policy optimization. The process of VAPO is shown in Fig 3.
Figure 3: Overview of the Visually-Anchored Policy Optimization framework. The OLLM takes audio, slide, and instruction as input, generates a structured output, and is optimized via reward functions that guide the policy update.
4.1 Structured Reasoning Format
Inspired by CoT Reasoning Jaech et al. (2024), we design a mandatory output format to guide the model in following the reasoning principle we call "Look before Transcription". Specifically, the model is required to generate its output within a unified structure of <think><answer>.
In the <think> block, the model first processes and digests the visual information. Before generating the final transcription, it must output recognized text from the slide image $I$. This step acts as an explicit internal OCR, ensuring the model sees and understands key slide information before processing the audio. The <answer> block is where the final transcription $Y$ is generated. Instead of generating directly, the model is required to reference the visual content anchored in the <think> block. This allows specialized terms and technical jargon from the slide to serve as prior knowledge, helping the model resolve ambiguous audio and homophones, improving the transcription of key named entities. Through this format, we decompose the complex and implicit SlideASR task into an ordered, two-stage explicit reasoning process.
4.2 Reinforcement Learning Optimization
We design four reward functions to guide the model’s learning comprehensively from different dimensions. To optimize the model’s policy with these rewards, we employ the Generative Representational Policy Optimization (GRPO) Shao et al. (2024b) algorithm for fine-tuning the model.
@@ 4.2 Reinforcement Learning Optimization
Format Reward. This reward aims to ensure that the model’s output strictly adheres to the <think></think><answer></answer> format. A positive reward is given if the model generates the complete structure. The reward function is as follows:
$R_{\text{Format}}=\begin{cases}1,&\text{If the format is correct}\\ 0,&\text{otherwise.}\end{cases}$ (2)
OCR Reward. This reward is used to evaluate the accuracy of the content in the <think> block. We compare the text generated by the model in <think> with the actual text on the slide. We calculate the reward using the Word Error Rate (WER). If the sample is in Chinese, each character is treated as a word. To ensure the reward is non-negative, we apply a clipping mechanism that prevents it from going below zero. We denote the text in the <think> block as $T_{t}$, and the text on the slide as $T_{s}$. The reward function is as follows:
$R_{\text{OCR}}=max(1-\text{WER}(T_{t},T_{s}),0)$ (3)
ASR Reward. This reward is used to evaluate the overall quality of the final transcription text in the <answer> block. We compare the output in <answer> with the ground truth speech transcription. Similarly, we apply the same clipping mechanism in ASR reward. We denote the text in the <answer> as $T_{a}$ block, and the ground truth speech transcription as $T_{g}$. The reward function is as follows:
$R_{\text{ASR}}=max(1-\text{WER}(T_{a},T_{g}),0)$ (4)
Visual Anchoring Reward. This reward in the VAPO is responsible for establishing a connection between the <think> and <answer> blocks. This reward specifically incentivizes the model to correctly use the named entities recognized in the <think> block within the <answer> block. The calculation process is as follows. First, we extract the set of correctly identified entities $E_{\text{think}}$ from the output of the <think>. Then, we calculate the F1 score of this entity set $E_{\text{think}}$ in the output of the <answer>. The reward function is as follows:
$R_{\text{VA}}=F1_{score}(E_{think},T_{answer}).$ (5)
This reward encourages the model to reference and utilize the slide information, effectively mitigating the issue of disjoint processing between the "seeing" and "hearing".
Finally, the model’s total reward is defined as a weighted sum of the four rewards described above:
$\begin{split}R_{total}=&\lambda_{1}R_{\text{Format}}+\lambda_{2}R_{\text{OCR}}\\ &+\lambda_{3}R_{\text{ASR}}+\lambda_{4}R_{\text{VA}}\end{split}$ (6)
$\lambda$ denotes the weighting hyperparameters for each individual reward.

### v2

# VAPO: End-to-end Slide-Enhanced Speech Recognition with Omni-modal Large Language Models

@@ Abstract
Omni-modal large language models (OLLMs) offer a promising end-to-end solution for slide-enhanced speech recognition due to their inherent multimodal capabilities. However, we found a fundamental issue faced by OLLMs: Visual Interference, where models show a bias towards visible text over auditory signals, causing them to hallucinate slide content that was never spoken. To address this, we propose Visually-Anchored Policy Optimization (VAPO), which aims to reshape models’ inference process to follow the human-like “Look-then-Listen” inference chain. Specifically, we design a temporally decoupled policy: the model first extracts visual priors in a <think> block to serve as semantic anchors, then generates the transcription in an <answer> block. The policy is optimized via multi-objective reinforcement learning. Furthermore, we introduce SlideASR-Bench, a comprehensive benchmark designed to address the scarcity of entity-rich data, comprising a large-scale synthetic corpus for training and a challenging real-world test set for evaluation. We conduct extensive evaluations demonstrating that VAPO effectively eliminates visual interference and achieves state-of-the-art performance on SlideASR-Bench and public datasets, significantly reducing entity recognition errors in specialized domains.
VAPO: End-to-end Slide-Enhanced Speech Recognition with Omni-modal Large Language Models
Rui Hu1,2, Delai Qiu2, Yining Wang2, Shengping Liu2, Jitao Sang1,3thanks: Corresponding author. 1Beijing Key Laboratory of Traffic Data Mining and Embodied Intelligence, Beijing Jiaotong University 2Unisound AI Technology Co., Ltd. 3State Key Laboratory of AI Safety, Beijing {rui.hu, jtsang}@bjtu.edu.cn https://github.com/isruihu/SlideASR-Bench
@@ 1 Introduction
Figure 1: Comparison of paradigms for the SlideASR task. (a) Pipeline paradigm: Cascades independent modules which is complex. (b) End-to-end paradigm: Illustrates two approaches using an OLLM. The Vanilla path directly generates the ASR result, often leading to visual interference. In contrast, our Look-then-Listen path enforces a structured inference chain, explicitly decoupling visual perception from auditory processing.
@@
Current Automatic Speech Recognition (ASR) models, such as Whisper Radford et al. (2023), have demonstrated impressive performance in general domains. However, recognition accuracy often deteriorates significantly in specialized scenarios, such as academic lectures or technical presentations where domain-specific terminologies and rare entities are prevalent Sinhamahapatra and Niehues (2025). While previous works have improved ASR accuracy by incorporating lip movement information Afouras et al. (2022); Ma et al. (2021, 2023); Shi et al. (2022), they ignore the rich semantic context present in presentation slides. In these scenarios, critical keywords and proper nouns are often explicitly displayed on the slides, serving as strong visual cues for the spoken content Wang et al. (2024a, b). For clarity, we refer to the task of improving ASR accuracy by incorporating visual context from presentation slides as SlideASR.
Currently, the dominant strategy for the SlideASR task is the pipeline paradigm Wang et al. (2024b). As shown in Fig. 1(a), this paradigm extracts text from slide images and selects domain-specific words, which are then fed as textual context into Large Audio Language Models (LALMs) Chu et al. (2024); Bai et al. (2024); Dinkel et al. (2025) for ASR. This cascading process involves multiple modules, suffering from complexity and error accumulation.
Given these limitations, a natural question arises: Can we establish an end-to-end paradigm that directly utilizes the slide image as visual context, thus bypassing the complex pipeline? To answer this question, we argue that the recently emerging Omni-modal Large Language Models (OLLMs) Xu et al. (2025a, b); Yao et al. (2024); Li et al. (2025a) offer a promising solution. OLLMs are capable of simultaneously processing textual, visual, and auditory modalities. Thus, they are inherently well-suited to accomplish the SlideASR task in an end-to-end manner.
However, our preliminary investigation reveals a critical gap between this theoretical potential and practical performance (Sec. 2). Specifically, we observe that current OLLMs suffer from severe visual interference. Instead of using the slide image as auxiliary information, the models exhibit a strong tendency to hallucinate, i.e., incorrectly transcribe visible slide text while this text is absent from the spoken audio. We attribute this failure to the lack of an intermediate reasoning phase. Considering human behavior, when listening to professional presentations, humans typically adopt a “Look-then-Listen” strategy, which temporally decouples the processing of the two modalities. We first scan the slide to establish a contextual prior of the topic and then anchor the auditory input within this context. In contrast, vanilla OLLMs process visual and auditory signals simultaneously (left path in Fig. 1(b)). Lacking this sequential guidance, strong visual cues tend to suppress auditory inputs, resulting in the observed interference.
Inspired by this, we propose Visually-Anchored Policy Optimization (VAPO) which centers on reshaping the model’s inference process to align with the human-like “Look-then-Listen” workflow (right path in Fig. 1(b)). Specifically, we design a <think><answer> format to explicitly decompose the multimodal task. In the <think> phase, the model is required to first perform Optical Character Recognition (OCR) to establish a visual context prior. Subsequently, in the <answer> phase, it generates the transcription by attending to the audio while referencing the extracted content as a reliable anchor. To instantiate this policy, we employ four complementary rewards: a) Format Reward to ensure structural compliance; b) OCR Reward to promote precise visual perception; c) ASR Reward to maintain the overall transcription accuracy; and d) Visual Anchoring Reward to encourage the model to effectively leverage key entities identified within the <think> phase.
Furthermore, to address data bottlenecks, we construct SlideASR-Bench. Existing datasets, e.g., SlideSpeech Wang et al. (2024b), primarily focus on general scenarios and lack sufficient entity density. SlideASR-Bench is specifically tailored to promote the utilization of visual cues for specialized terms and comprises two subsets: 1) SlideASR-S, a large-scale synthetic corpus derived from ContextASR-Bench Wang et al. (2025) providing both training and test sets; and 2) SlideASR-R, a small real-world test set for evaluation in complex presentation environments.
@@
We conduct extensive experiments on SlideSpeech Wang et al. (2024b) and our proposed SlideASR-Bench. Empirical results demonstrate that our approach significantly outperforms state-of-the-art models, e.g., Qwen3-Omni Xu et al. (2025b), particularly on entity-related metrics. The main contributions are summarized as follows:
  • Analysis: We identify visual interference as the primary bottleneck in current OLLMs for SlideASR, revealing why naive end-to-end approaches fail.
  • Method: We propose VAPO, which reshapes the inference process into a structured, human-like “Look-then-Listen” workflow, explicitly temporally decoupling visual perception from auditory transcription to eliminate visual interference.
  • Data: We construct SlideASR-Bench, comprising synthetic and real-world subsets, to train and evaluate SlideASR task in entity-rich scenarios.
2 Visual Interference of OLLMs in End-to-end SlideASR
Although OLLMs theoretically have the capability to process auditory and visual modalities simultaneously, our preliminary investigation reveals a critical failure mode in their practical application. We present a motivating example in Fig. 2 using Qwen2.5-Omni-7B Xu et al. (2025a). In the audio-only setting, the model accurately transcribes the speech. However, when the corresponding slide image is introduced, the model’s behavior shifts drastically: instead of utilizing the visual information as auxiliary context, it ignores the auditory signal and directly outputs the text visible on the slide. We term this specific failure mode as Visual Interference. Formally, it is defined as a phenomenon where the model generates text present in the visual context while it is absent from the speech.
Figure 2: An illustrative example of Visual Interference in the end-to-end SlideASR task. Left: the model correctly transcribes the speech in audio-only mode. Right: the model erroneously outputs the text on the slide when the image is introduced as visual context.
To systematically quantify this phenomenon, we introduce a metric termed the Visual Interference Rate (VIR) which measures the tendency of the model to hallucinate text from the slide that was never spoken. We evaluate four representative OLLMs: Qwen2.5-Omni (7B/3B) Xu et al. (2025a), MiniCPM-o-2.6 Yao et al. (2024), and Megrez-Omni Li et al. (2025a) on the SlideSpeech Wang et al. (2024b) dataset. The calculation follows a set-difference logic:
  • Step 1: Context Isolation. Let $V_{slide}$ be the set of words on the slide and $V_{audio}$ be the set of words in the ground-truth transcript. We first identify the slide-exclusive vocabulary $V_{exclusive}=V_{slide}\setminus V_{audio}$. These are words that appear visually but are not spoken.
  • Step 2: Interference Detection. Let $V_{pred}$ be the set of words in the model’s prediction. We calculate the intersection $I=V_{pred}\cap V_{exclusive}$.
  • Step 3: Metric Calculation. If $I\neq\emptyset$, the sample is flagged as exhibiting visual interference. The VIR is defined as the percentage of such flagged samples in the dataset.
The quantitative results are presented in Table 1. All evaluated OLLMs exhibit a high VIR, ranging from 12.87% to as high as 63.28% on the test set. Even the strongest baseline, Qwen2.5-Omni-7B, fails to suppress visual interference in over 12% of the samples. This consistent failure across different models indicates that this is a widespread issue rather than an isolated case. This phenomenon highlights a fundamental limitation in current paradigms: OLLMs lack an explicit mechanism to temporally decouple visual perception from auditory processing.
Table 1: Comparison of Visual Interference Rate across different OLLMs on the SlideSpeech dataset.
Model Dev set Test set
@@ 2 Visual Interference of OLLMs in End-to-end SlideASR
3 Method
Figure 3: Overview of the Visually-Anchored Policy Optimization (VAPO) framework. The model takes audio, slide, and instruction as inputs and generates a structured <think><answer> sequence. In the <think> phase, the model extracts visual context, which serves as a semantic anchor (indicated by the red box) to guide the transcription in the <answer> phase. The policy is optimized via four rewards (Format, OCR, ASR, and Visual Anchoring).
To mitigate the Visual Interference issue identified in Sec. 2, we propose Visually-Anchored Policy Optimization (VAPO). Inspired by the human perception process, VAPO addresses this limitation by reshaping the model’s inference process to achieve a temporal decoupling of modalities. Specifically, it establishes a structured <think><answer> inference chain: the model first establishes a visual prior in the <think> phase, and then references this content as an anchor to guide speech transcription in the <answer> phase. This mechanism mimics the human-like “Look-then-Listen” workflow. The overall framework is illustrated in Fig. 3.
3.1 The “Look-then-Listen” Inference Chain
We design a structured output format to explicitly enforce the necessary temporal decoupling between visual perception and auditory processing. Drawing inspiration from Chain-of-Thought (CoT) reasoning Jaech et al. (2024); Ma et al. (2025), the model is mandated to generate its output within a sequential <think><answer> structure.
The <think> block corresponds to the “Look” phase: the model first processes the visual input and is tasked with extracting textual information from the slide image. This operation establishes a critical visual context prior before the auditory processing begins. Subsequently, the <answer> block corresponds to the “Listen” phase: the model generates the final transcription. In this phase, the model is required to reference the content anchored in the <think> block. This mechanism allows specialized terms from the slide to serve as semantic anchors, thereby assisting the model in resolving ambiguous audio and enhancing overall transcription accuracy.
3.2 Multi-Objective Policy Optimization
To make the model follow the proposed inference chain, we design four complementary reward functions to guide the model’s learning. We employ the Group Relative Policy Optimization (GRPO) algorithm Shao et al. (2024b) for model training.
@@ 3.2 Multi-Objective Policy Optimization
Format Reward. This reward aims to ensure that the model’s output strictly adheres to the <think><answer> format. A positive reward is assigned only if the model generates the complete structural tags. The reward function is defined as:
$R_{\text{Format}}=\begin{cases}1,&\text{if the format is correct,}\\ 0,&\text{otherwise.}\end{cases}$ (1)
OCR Reward. This reward evaluates the visual perception quality in the “Look” phase by comparing the text generated in the <think> block ($T_{t}$) with the ground truth slide text ($T_{s}$). We use the Word Error Rate (WER) as the metric, treating each Chinese character as a word where applicable. The reward is normalized and clipped to ensure non-negativity:
$R_{\text{OCR}}=\max(1-\text{WER}(T_{t},T_{s}),0).$ (2)
ASR Reward. This reward assesses the transcription quality in the “Listen” phase by comparing the output in the <answer> block ($T_{a}$) with the ground truth speech transcription ($T_{g}$). Similarly, it is derived from the WER:
$R_{\text{ASR}}=\max(1-\text{WER}(T_{a},T_{g}),0).$ (3)
Visual Anchoring Reward. This reward bridges the “Look” and “Listen” phases by incentivizing the model to ground spoken entities in the visual context. Let $E_{target}$ be the set of critical entities present in both the ground truth slide and the ground truth speech transcript. Let $E_{gen}$ be the subset of $E_{target}$ that successfully appears in both the generated <think> and <answer> blocks. The reward is defined as recall of these target anchors:
$R_{\text{VA}}=\begin{cases}\frac{|E_{gen}|}{|E_{target}|},&\text{if }|E_{target}|>0,\\ 0,&\text{otherwise.}\end{cases}$ (4)
This formulation encourages the model to explicitly capture and utilize the visual cues that are relevant to the current speech.
Finally, the total reward is defined as a weighted sum of the four components, where $\lambda_{i}$ denotes the weighting hyperparameter for each reward:
$\begin{split}R_{\text{total}}=\;&\lambda_{1}R_{\text{Format}}+\lambda_{2}R_{\text{OCR}}\\ &+\lambda_{3}R_{\text{ASR}}+\lambda_{4}R_{\text{VA}}.\end{split}$ (5)